# GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1

### v1

@@ IV-B Integrated Mapping
The second main contribution of this work is an integrated mapping algorithm that incorporates the mapping directly into the graph partitioning. We follow the multilevel approach, but instead of minimizing the edge-cut we optimize $J(C,D,\Pi)$. The multilevel approach, coarsens the graph to a small representation, determines an initial partition and afterwards uncoarsens the graph and refines the mapping. We will detail our methodology for each phase in the next few paragraphs. We closely follow the matching, coarsening, refinement and rebalancing methods used by the GPU-based graph partitioner Jet [18] and adapt them to process mapping.
Edge-List Graph Format. Before presenting the multilevel process mapping scheme, we want to present our extended CSR format. As introduced in Section III-D, the CSR representation naturally supports parallelization over vertices, allowing each thread to process a vertex’s adjacency list independently. However, some computations benefit from direct parallelization over edges. In standard CSR, this requires nested parallel loops, first over vertices, then their neighbors, which leads to poor load balancing due to varying vertex degrees, resulting in idle threads on GPUs. To overcome this, we extend the CSR format with an additional array $E_{u}$ of size $|E|$ that explicitly stores the opposite endpoint of each edge. This edge-centric representation enables flat parallelism over all edges, improving load balance and GPU utilization. We refer to this edge-list representation as $\mathbb{E}$ throughout the rest of this work.
@@
Matching. We utilize the two-hop matching presented by LaSalle et. al [29] but make a few changes. It starts with a heavy-edge matching algorithm. Instead of using $\omega(\{u,v\})$ as the weight of the edge, we use the rating function $\exp^{*2}(\{u,v\})=\frac{\omega(\{u,v\})^{2}}{c(u)c(v)}$ presented by Holtgrewe et al. [26] and also used in IntegratedMapping [15]. Their experiments showed that it lead to improved running times and solution quality in the context of graph partitioning. By looping over $\mathbb{E}$ in parallel, we determine for each vertex $v$ its preferred neighbor $p(v)=u$ as the one with the greatest rating. Adjacent vertices $u$ with $c(v)+c(u)>L_{\max}$ are excluded to prevent an imbalanced initial solution. Afterwards, we determine in parallel for each vertex $v$ if $p(p(v))=v$, i.e., vertex $v$ prefers $p(v)$ and $p(v)$ prefers $v$. These vertices are matched to each other. If the vertices have multiple neighbors with the same rating, then it is very unlikely that $v$ and $p(v)$ prefer each other, and we would end up with few matches. We therefore add a small amount of deterministic noise $\eta(\{u,v\})$ onto each rating. The noise breaks the equality and is small enough to not influence the order of two distinct ratings. The matching procedure can be repeated multiple rounds, excluding previously matched vertices. Afterwards, leaf, twin and relative matchings are also applied multiple times if necessary. We adopt the same hashing and matchmaker vertex strategy of Jet to improve performance. Instead of matching at least 75% of vertices, we reduce the percentage to 40%. While this leads to more levels in the multilevel framework, it also allows for more chances to refine the partition in the refinement phase. In our preliminary experiments using greater values lead to drastically worse solution quality.
@@
Coarsening. Edge-contraction, to create the coarsened graph, is based on a fine-grained per-vertex hashing scheme from Jet [18]. The pseudocode can be seen in Algorithm 2. Let $G=(V,E)$ be the current graph and let $M:V\mapsto[n_{c}]$ map each vertex to its coarse counterpart. The value $n_{c}$ is the number of vertices in the coarse graph. The algorithm first computes an upper bound on the neighborhood size of each coarse vertex by summing the degrees of all vertices assigned to it (Lines 131-3). In Line 4 and 5 two hash arrays, $H_{v}$ and $H_{w}$, are allocated that will hold the edges and their weight. The offset array $O$ denotes the index range $O[v]$ to $O[v+1]$ where the edges of the coarse vertex $v$ are placed. Next all edges $(u,v,w)\in\mathbb{E}$ are processed in parallel. If both endpoints contract to the same vertex, the edge becomes a self loop and is discarded (Line 8). Otherwise we need to insert $(M(v),w)$ into the adjacency of $M(u)$ in the interval $\mathcal{I}$. Note that adding $(M(u),w)$ into the adjacency of $M(v)$ is handled by the work unit with reverse edge $(v,u,w)\in\mathbb{E}$. Insertion is done in Line 10, which we will explain in the next paragraph. After the parallel loop ends it remains to extract the actual graph from the arrays $H_{v}$ and $H_{w}$. This boils down to determining the true neighborhood sizes of each vertex and extracting the edges by skipping all Null entries in $H_{v}$.
The most important part of our fine grained edge-contraction is the InsertOrLookup function in Line 10. To handle concurrent insertions without locks we utilize atomic compare-and-swap (CAS) operations. A CAS operation compares the value at $H_{v}[j]$ with Null and only if so swaps $M(v)$ in its place. Regardless if the check succeeded the previous value is returned. This means, if a thread executes a CAS operation on $H_{v}[j]$ and Null or $M(v)$ is returned, a slot for $M(v)$ was found. The weight $w$ will be atomically added to $H_{w}[j]$. Otherwise the thread will continue to cycle through the interval $\mathcal{I}$. Since we overestimated the neighborhood size of $M(u)$ at least one open slot will be available. To reduce atomic contention on $H_{v}[j]$, each thread determines $j\in\mathcal{I}$ by hashing $M(v)$ and then cycling through the range if necessary.
Algorithm 2 GPU-based Coarsening
@@
10:  $\smash{\mathcal{I}\leftarrow[O[M(u)],O[M(u)+1])}$
11:  $\smash{\textsc{InsertOrLookup}(H_{v},H_{w},\mathcal{I},M(v))}$
12:return $\smash{\textsc{ExtractCSR}(B,H_{v},H_{w})}$
@@
Initial Partitioning. Initial partitioning is performed once the last coarsened graph has fewer than $128k$ vertices. Choosing a smaller threshold often resulted in an imbalanced initial partitions from which the algorithm failed to recover. We apply the hierarchical multisection approach explained in Section IV-A to acquire an initial mapping, but execute it entirely on the CPU. At this stage, the graph is sufficiently small that GPU execution offers no runtime advantage. Moreover, even the GPU-based partitioner Jet [18] would internally quickly fall back to the CPU-based partitioner Metis [27] when presented with such small graphs. To execute hierarchical multisection, we utilize the CPU-based partitioner KaFFPa [40] in its Fast configuration. Although the Strong configuration yields higher-quality mappings, its significantly increased runtime would dominate the total execution time, rendering the overall algorithm non-competitive.
@@
Refinement. We adapt the unconstrained label propagation of Jet [18] to use $J(C,D,\Pi)$ as the objective function. The corresponding pseudocode is provided in Algorithm 3. The refinement algorithm begins by determining the best target block $\Pi^{\prime}(v)$ for each vertex $v$ by maximizing the gain $G_{b}(v)$ (see Equation 1) across all neighboring blocks (Lines 353-5).
@@
1:Graph $\smash{G}$, Mapping $\Pi$, Is-Locked array $L$
@@
4:for $v\in V$ do in parallel
5:  if $\smash{L[v]=1}$ then skip $\triangleright$ Vertex locked
6:  $\smash{B\leftarrow\{\Pi(u)\mid u\in N(v)\}}$
7:  $\smash{\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)}$
8:  if $\smash{G_{\Pi^{\prime}(v)}(v)\geq 0}$ then $\smash{\mathcal{X}[v]=1}$
@@
10:for $v\in V$ do in parallel
11:  if $\smash{\mathcal{X}[v]=1\text{ and }\smash{\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0}}$ then $\smash{\mathcal{M}[v]=1}$
@@
13:return $\smash{\mathcal{M},\Pi^{\prime},L}$
@@
In Jet, a filter is applied that retains only vertex moves satisfying $G_{b}(v)\geq 0\lor-G_{b}(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$. This criterion is suitable when minimizing the edge-cut, as both $G_{b}(v)$ and $\operatorname{conn}(v,\Pi(v))$ depend solely on edge weights and are thus directly comparable. In contrast, for process mapping, $G_{b}(v)$ contains the additional communication factors and is therefore often much smaller than $\operatorname{conn}(v,\Pi(v))$ and varies significantly, depending on $v$’s connections. This renders the original condition ineffective. We therefore only allow non-negative moves $G_{b}(v)\geq 0$ to pass the filter, as it performed best in our preliminary experiments. A more sophisticated filtering criterion may further improve solution quality and remains a topic for future work.
In Line 7, vertices passing this filter are inserted into the set $\mathcal{X}$, which is implemented as a binary array marking selected vertices. The set is implicitly ordered by $\operatorname{ord}$, defined as
@@
$\operatorname{ord}(u)\begin{cases}<\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G(u)>G(v),\\ <\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G(u)=G(v)\land u<v,\\ >\operatorname{ord}(v)&\text{otherwise}.\end{cases}$
@@
Next, all vertex moves in $\mathcal{X}$ are reevaluated under an approximated future mapping state. If a vertex $v$ has a neighbor $u$ scheduled to move to block $\Pi^{\prime}(u)$ with $\operatorname{ord}(u)<\operatorname{ord}(v)$, then $u$ is assumed to have already moved when recomputing the gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$. We mark each vertex $v\in\mathcal{X}$ in a To-Move array $\mathcal{M}$ if it has non-negative gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ (Lines 9109-10). To prevent oscillation of vertices, the To-Move array serves as the locked array $L$ for the next iteration (Line $11$). The function returns the To-Move array $\mathcal{M}$ the destination mapping $\Pi^{\prime}$ and the Locked array $L$.
Rebalancing. Our rebalancing is also adapted from Jet [18]. Jet splits its rebalancing into a weak and strong algorithm. The weak rebalancing keeps the loss small, however, it might need multiple iterations to achieve a balanced partition. The strong rebalancing instead often achieves a balanced partition in one iteration, but has a greater loss.
@@
Weak rebalancing starts by determining for each vertex $v$ in an overloaded block, the move with maximum gain (or equally minimum loss) $G_{b}(v)$ into a neighboring block $b$ that has a weight smaller than the threshold $\sigma\leq L_{\max}$. The threshold $\sigma$ is chosen such that there is a small dead-zone between $\sigma$ and $L_{\max}$, so $b$ can take multiple vertices without instantly being overloaded. In our implementation we start with $\sigma=L_{\max}-0.065L_{\max}$ and gradually lower it throughout the multilevel scheme until at the lowest level we use $\sigma=L_{\max}-0.005L_{\max}$. We lower it, since the vertex weights are also gradually getting smaller during uncoarsening. Therefore, we can operate on a more fine level and more blocks become eligible. If for a vertex no such neighboring block exists, a random block that fits the threshold is chosen.
@@
Ideally, each block would store all its outgoing vertex moves decreasingly in a list and then a prefix of the list could be chosen, such that the block becomes balanced. This would grant the smallest possible loss based on the current partition. However, since sorting can be expensive on a GPU, we utilize an approximation of the sorted list similar to Jet. Each vertex move is assigned a slot in the list $0,-1,\ldots,-10,-20,\ldots,-100,-200,\ldots,-1000,\operatorname{X}]$. The slot ++ holds all positive gain moves, the slot $\operatorname{X}$ contains all very large negative gain moves and all other slots hold moves with gain $L[i]\geq G_{b}(v)>L[i+1]$. Note that at first we increase the range by 1, then by 10 and in the end by 100. This gives us an increased resolution for moves with a comparatively small loss, which minimizes the loss when only a few vertices need to be moved. Note that Jet also adopts the ++ and 0 slot and for any negative moves they use $\log_{2}(-G_{b}(v))$ to assign a vertex a slot. The approximated list is build in four phases, first by counting how many vertices fall into each slot, allocating enough memory, building a exclusive prefix-sum over the sizes and in the last phase inserting the vertices into their slots. To reduce atomic contention we utilize $\rho=2$ mini buckets in each slot and a vertex $v$ is assigned its mini bucket by $v\mod\rho$. As the last step of weak rebalancing a list prefix is chosen for each block $b$ and all vertices in it are moved. The prefix moves at least $c(b)-L_{\max}$ weight, so that the block $b$ becomes balanced.
@@
Since the lists are approximately sorted by their gain, the prefixes have a small combined loss. However, there is no safeguard against the destination blocks becoming overloaded. While the threshold $\sigma$ helps in preventing this, it does not guarantee it. Therefore, the destination blocks might become overbalanced and another iteration is required.
Strong rebalancing operates in a similar manner to weak rebalancing but reverses the assignment perspective. As before, each vertex $v$ in an overloaded block $b$ identifies an eligible neighboring target block $b^{\prime}$ that minimizes the loss of moving $v$. However, instead of storing the vertex moves with the source block $b$, they are stored in the lists of the corresponding target block $b^{\prime}$. Each target block then selects a prefix of moves from its own list, such that it stays balanced after accepting them. After execution, the target blocks are guaranteed to stay balanced, while the source blocks lose weight and are also likely to be balanced. Consequently, strong rebalancing can restore balance in just one iteration. However, target blocks will often move more vertices than necessary, resulting in a higher overall loss compared to weak rebalancing.
@@
7:   $\mathcal{M},$ $\Pi^{\prime\prime}$, $L$ $\leftarrow\operatorname{UnconLabelProp}(G,\Pi^{\prime},L)$

### v2

@@ 4.2 Integrated Mapping
The second main contribution of this work is a GPU-based integrated mapping algorithm that incorporates the mapping directly into the graph partitioning. We follow the multilevel approach, but instead of minimizing the edge-cut, we minimize $J(\mathcal{C},\mathcal{D},\Pi)$. Remember that the multilevel approach first coarsens the graph into a sequence of progressively smaller graphs. The smallest graph is then partitioned and the coarsening is reverted. Refinement algorithms are used during uncoarsening to further optimize the objective. We will detail our methodology for each phase in the next few paragraphs. We closely follow the matching, coarsening, refinement and rebalancing methods used by the GPU-based graph partitioner Jet [19] and adapt them to process mapping.
@@
Matching. To determine a matching, the two-hop matching presented by LaSalle et al. [30] is adapted with a few minor changes. It starts with a heavy-edge matching algorithm. Instead of using $\omega(\{u,v\})$ as the rating of the edge, the rating function $\exp^{*2}(\{u,v\})=\frac{\omega(\{u,v\})^{2}}{c(u)c(v)}$ presented by Holtgrewe et al. [27] is used. Their experiments show that it leads to improved running times and solution quality in the context of graph partitioning. In parallel, each thread processes one vertex $v$. All unmatched neighbors are scanned and the vertex $u$ with the highest rating is saved as the preferred neighbor $p(v)=u$. To prevent identical ratings, a small amount of deterministic noise $\eta(\{u,v\})$ is added to each rating. It breaks any equalities, but is small enough to not influence the result. Vertices $u$ with $c(v)+c(u)>L_{\max}$ are excluded to prevent an imbalanced initial solution. Afterwards, each vertex $v$ is again processed in parallel. If $p(p(v))=v$, i.e., vertex $v$ prefers $p(v)$ and $p(v)$ prefers $v$, then $v$ and $p(v)$ are matched to each other. The matching procedure is repeated for multiple rounds until no more matches are determined. If more than $75\%$ of vertices are matched then the matching stops. Else, leaf, twin and relative matchings are applied to increase the number of matched vertices. We adopt the same hashing and matchmaker vertex strategy of Jet to improve performance. In process mapping, matching-based coarsening approaches are usually advantageous, since the task graphs stem from scientific simulations and have a mesh-like structure. Another approach would be to cluster the graph, however this is usually superior in unstructured workloads like social networks.
@@
Contraction. Edge contraction, to create the coarse graph, is based on a fine-grained per-vertex hashing scheme from Jet [19], however, it is extended by using the extended CSR format. The overall idea is to assign each coarser vertex a hash array and the threads then hash the coarser edges into the respective array. The pseudocode can be seen in Algorithm 3. Let $G=(V,E)$ be the current graph, $n_{c}$ the number of vertices in the coarse graph and $M:V\mapsto[n_{c}]$ map each vertex to its coarse counterpart. The algorithm first computes an upper bound on the neighborhood size of each coarse vertex by summing the degrees of all vertices assigned to it (Lines 131-3). In Line 4 and 5 two hash arrays, $H_{v}$ and $H_{w}$, are allocated that will hold the edges and their weights. The offset array $O$ denotes the index range $[O[v],O[v+1])$ where the edges of the coarse vertex $v$ are placed. Then, all edges $(u,v,w)\in\mathbb{E}$ are processed in parallel. If both endpoints contract to the same vertex, the edge becomes a self-loop and is discarded (Line 8). Otherwise, we need to insert $(M(v),w)$ into the adjacency of $M(u)$ in the interval $\mathcal{I}$. Note that adding $(M(u),w)$ into the adjacency of $M(v)$ is handled by the work unit with reverse edge $(v,u,w)\in\mathbb{E}$. Insertion is done in Line 10, which we will explain in the next paragraph. After the parallel loop ends, it remains to extract the actual graph from the arrays $H_{v}$ and $H_{w}$. This boils down to determining the true neighborhood sizes for all vertices, building the offset and extracting the edges by skipping all Null entries in $H_{v}$. This can be accomplished by two parallel_scans.
Algorithm 3 GPU-based Contraction
@@
10:  $\mathcal{I}\leftarrow[O[M(u)],O[M(u)+1])$}
11:  $\textsc{InsertOrLookup}(H_{v},H_{w},\mathcal{I},M(v))$}
12:return $\textsc{ExtractCSR}(B,H_{v},H_{w})$}
The most important part of the contraction is the InsertOrLookup function in Line 10. Each thread must place its coarse edge $(M(v),w)$ into $H_{v}$ and $H_{w}$ in the interval $\mathcal{I}$. In a serial algorithm one would first check if the edge $(M(v),w)$ already has an entry in $H_{v}$ and if so, add the weight to it. If there is no such entry then one would place the edge at a Null entry. However, in a parallel setting multiple threads add edges simultaneously and it must be ensured that edges are not overwritten and that duplicate edges are summed properly. To handle concurrent insertions, atomic compare-and-swap (CAS) operations can be utilized. A CAS operation $\text{CAS}(\textsc{ptr},\textsc{expected},\textsc{desired})$ compares the value at memory address ptr, with the expected vale and only if equal, writes the desired value in its place. Regardless of whether the thread wrote desired or not, the previous value stored at ptr is always returned. The important part of the CAS operation is that the check and the write are executed as one atomic operation. Two threads can never read the same expected value and write different desired values.
A thread now inserts its edge by executing a CAS operation on memory address $H_{v}[j]$ with the expected value of Null and the desired value $M(v)$. If $M(v)$ is returned, then the edge was already written in $H_{v}[j]$ and the weight $w$ is atomically added to $H_{w}[j]$. If Null is returned, then this thread just wrote $M(v)$ into $H_{v}[j]$ and $w$ is also atomically added to $H_{w}[j]$. If neither is returned, then $H_{v}[j]$ already holds a different edge. The thread will continue cycling through the interval $\mathcal{I}$ until a valid position has been found. Since the neighborhood size of $M(u)$ was overestimated, at least one open slot will be available and the thread is able to place its edge. To reduce atomic contention on $H_{v}[j]$, each thread determines $j\in\mathcal{I}$ by hashing $M(v)$ and then cycling through the interval if necessary.
@@
Initial Partitioning. Initial partitioning is performed once the last coarsened graph has fewer than $8k$ vertices. We apply the hierarchical multisection [48, 45] approach, explained in Section 4.1, to acquire an initial mapping, but execute it entirely on the CPU. At this stage, the graph is sufficiently small that GPU execution offers no runtime advantage. Moreover, even a GPU-based graph partitioner like Jet would internally quickly fall back to a CPU-based partitioner when presented with such small graphs. Also currently no GPU-based initial partitioning algorithms exist for HPMP or GP. To execute hierarchical multisection, we utilize a simple $k$-way graph partitioner.
Uncontraction. Each coarser graph $G,G_{1},\cdots,G_{t}$ is separately stored in memory so uncontracting the graph from $G_{i+1}$ to $G_{i}$ is achieved by simply deleting $G_{i+1}$. The mapping $\Pi$, however, has to be updated. Let $M$ be the mapping that was used to create the coarser graph $G_{i+1}$ from $G_{i}$. Then set in parallel $\Pi(v)=\Pi(M(v))$ for each $v\in G_{i}$.
Refinement. We adapt the unconstrained label propagation of Jet [19] to use $J(\mathcal{C},\mathcal{D},\Pi)$ as the objective function instead of edge-cut. Classic label propagation visits all vertices in a random order and if moving the current vertex to a neighboring block improves the objective and adheres to the balance constraint, then the move is made immediately. Therefore, the objective value never decreases and the balance constraint is always fulfilled. In a parallel setting, one would like to determine for each vertex in parallel whether moving it is beneficial. This part can easily be parallelized. Then, the beneficial moves would ideally also be executed in parallel. However, when moving the vertices in parallel, difficulties arise. Each individual move might adhere to the balance constraint, but moving multiple vertices at once can quickly overload a block. Also, although each individual move might be beneficial, moving adjacent vertices in parallel can have an overall negative effect on the solution quality. This makes it inherently difficult to parallelize label propagation.
Jet’s approach works in two phases. At first an unconstrained label propagation algorithm is used that determines beneficial vertex moves. In this phase the balance constraint is completely ignored and subsequently a second rebalancing phase is required. This phase aims to restore the balance, while keeping the loss in the objective function small.
The unconstrained label propagation works in two phases. At first, for each vertex the best move to a neighboring block is determined. These moves are filtered and only non-negative moves are considered further. Each move is then reevaluated under a future approximate partition state. If the move stays positive, then it is finally saved to be executed.
@@
The pseudocode of our adapted refinement algorithm is provided in Algorithm 4. The refinement algorithm begins by determining the best target block $\Pi^{\prime}(v)$ for each vertex $v$ by maximizing the gain $G_{b}(v)$ (see Equation 1) across all neighboring blocks (Lines 4 – 6). If indeed a move with non-negative gain has been found, then we insert the vertex into the list $\mathcal{X}$ via an atomically incremented index (Line 7).
@@
1:Graph G{G}, Mapping $\Pi$
@@
5:for $v\in V$ do in parallel
6:  $B\leftarrow\{\Pi(u)\mid u\in N(v)\}$}
7:  $\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)$}
8:  if $G_{\Pi^{\prime}(v)}(v)\geq 0$} then $\mathcal{X}[i++]=v$}
@@
[…]
11:for $v\in\mathcal{X}$ do in parallel
12:  if $\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0$} then $\mathcal{M}[i++]=v$}
13:return $\mathcal{M},\Pi^{\prime}$}
@@
Originally, Jet applies a more relaxed filter that also retains negative vertex moves. A vertex $v$ must only satisfy $G_{b}(v)\geq 0\lor-G_{b}(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$. The value $\operatorname{conn}(v,\Pi(v))$ denotes the sum of edge weights connecting $v$ and block $\Pi(v)$. Negative vertex moves are allowed to pass this filter, such that the algorithm may escape local minima. The criterion is suitable when minimizing the edge-cut, as both $G_{b}(v)$ and $\operatorname{conn}(v,\Pi(v))$ depend solely on edge weights and are thus directly comparable. In contrast, for process mapping, $G_{b}(v)$ contains the additional communication factors and is therefore often much greater than $\operatorname{conn}(v,\Pi(v))$. This renders the original filter ineffective. Therefore only non-negative moves $G_{b}(v)\geq 0$ are allowed to pass the filter and be inserted into $\mathcal{X}$, as it performed best in our preliminary experiments. A more sophisticated filtering criterion may further improve solution quality and remains a topic for future work.
The list $\mathcal{X}$ now holds all non-negative vertex moves. In a serial algorithm, one could now sort the list and execute the best vertex moves first. Moves further to the end of the list could then be reevaluated under the new partition state and only be executed if they are positive. However, in a parallel setting this is not straightforward. Firstly, sorting on a GPU is expensive and should be avoided and secondly, executing one move after the other is inherently serial. Gilbert et al. [19] therefore first only implicitly sort the list $\mathcal{X}$ and then each vertex move is re-evaluated under an approximate future partition state in parallel. The list $\mathcal{X}$ is unordered, but we can define an ordering on its elements via
@@
$\operatorname{ord}(u)\begin{cases}<\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G_{\Pi^{\prime}(u)}(u)>G_{\Pi^{\prime}(v)}(v),\\ <\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G_{\Pi^{\prime}(u)}(u)=G_{\Pi^{\prime}(v)}(v)\land u<v,\\ >\operatorname{ord}(v)&\text{otherwise}.\end{cases}$
@@
Vertices with a high gain are at the front of the list, while vertices with lower gain are towards the end of the list. By querying $G_{\Pi^{\prime}(v)}(v)\geq 0$} we can determine whether a vertex was inserted into $\mathcal{X}$. It is now assumed that vertices with a lower ordering would have moved in a serial algorithm before a vertex with a higher ordering. All vertex moves in $\mathcal{X}$ are now reevaluated in parallel under an approximate future mapping state. We denote the reevaluated gain by $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ and it is determined just like the gain $G_{\Pi^{\prime}(v)}(v)$ (Equation 1), only the assigned block of some neighbors change. If a vertex $v$ has a neighbor $u$ with $\operatorname{ord}(u)<\operatorname{ord}(v)$ and vertex $u$ is scheduled to move to block $\Pi^{\prime}(u)$, then $u$ is assumed to have already moved when recomputing the gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$. Each vertex $v\in\mathcal{X}$ that has a non-negative gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ is then added to the final list of moves (Lines 9 – 11). To prevent oscillation of vertices, all vertices in the final list $\mathcal{M}$ are locked for the next iteration of label propagation. The results are also cached and if the neighborhood of a vertex did not change, then its result is reused. Both of these techniques are not shown in Algorithm 4 for visual clarity. The function returns the list $\mathcal{M}$ and the destination mapping $\Pi^{\prime}$.
Rebalancing. The rebalancing is also adapted from Jet [19], which splits its rebalancing into a weak and strong part. The idea is to determine for each vertex in an overloaded block the move with minimal loss to an underloaded block. All moves are approximately sorted and then as few vertices as necessary are moved. Weak rebalancing aims to keep the loss small, but might need multiple iterations, while strong rebalancing balances in one iteration, but has a much greater loss.
@@
The weak rebalancing starts by determining for each vertex $v$ in an overloaded block the move with maximum gain (or equally minimum loss) $G_{b}(v)$ into a neighboring block $B$ that has a weight less than the threshold $\sigma\leq L_{\max}$. The threshold $\sigma$ is chosen such that there is a small dead zone between $\sigma$ and $L_{\max}$, so $B$ can take multiple vertices without instantly being overloaded. In our implementation we set $\sigma=L_{\max}-100$, just like Jet. If for a vertex no such neighboring block exists, a random block that satisfies the threshold is chosen. Heavy vertices $v$ that weigh more than $1.5(c(\Pi(v))-c(V)/k)$ are not allowed to move during rebalancing.
@@
Ideally, each block would store all its outgoing vertex moves in decreasing order in a list and then a prefix of the list could be chosen, such that the block becomes balanced. This would grant a relatively small loss based on the current partition. However, sorting on a GPU is expensive, so instead an approximation of the sorted list is utilized. Each negative vertex move is assigned a slot in a $\log_{2}$–spaced bucket list. A vertex $v$ with negative gain $G_{b}(v)$ would therefore be stored in bucket $i$ with $i\leq\log_{2}(-G_{b}(v))<i+1$. This gives an increased resolution for moves with a comparatively small loss, which minimizes the loss when only a small number of vertices need to be moved. Strictly positive vertex moves are assigned to a special ++-bucket and 0-gain moves are placed in a special 0-bucket.
Instead of first building the approximation list explicitly and then determining the prefix of vertex moves, one can also use a per-vertex decision process. For each vertex, when it is added to a bucket it will store the accumulated weight of vertices already inside of its bucket. After all vertices have been processed, a prefix sum over all bucket weights is computed. Through this, each vertex can now determine the sum of the vertex weights in earlier buckets and it knows how much vertex weight will be moved in its own bucket, before the vertex itself is moved. If this weight is smaller than $c(B)-L_{\max}$, then this vertex needs to be moved.
Algorithm 5 Weak Rebalancing
1:Graph $G=(V,E)$, Mapping $\Pi$
[…]
3:for $v\in V$ do in parallel
4:  if $c(\Pi(v))\leq L_{\max}$ or $v$ too heavy do Skip
5:  $B\leftarrow\{\Pi(u)\mid u\in N(v)\land c(\Pi(u))\leq\sigma\}$}
6:  if $B=\emptyset$ then
7:   $B\leftarrow\{\text{rnd Block b with }c(b)\leq\sigma\}$
8:  $g\leftarrow\max_{b\in B}G_{b}(v)$
9:  $\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)$}
[…]
12:for $v\in V$ do in parallel
13:  if $\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0$} then $\mathcal{M}[i++]=v$}
14:return $\mathcal{M},\Pi^{\prime}$}
@@
By approximately sorting the lists by gain, and only moving as few vertices as possible, the overall combined loss stays small. However, there is no safeguard against the destination blocks becoming overloaded. While the threshold $\sigma$ helps in preventing this, it does not guarantee it. Therefore, the destination blocks might become overloaded and another iteration is required. The strong rebalancing operates in a very similar manner to weak rebalancing but takes extra precaution to not overload the destination blocks. Vertices that would overload a destination block are reassigned to other underloaded blocks. However, the vertices are not necessarily connected to their new destination block, which results in a much greater loss.
We adapted the method to use $J(\mathcal{C},\mathcal{D},\Pi)$ as the objective function. Although this increases computational complexity, we observed no improvement in solution quality. Therefore, we use the standard rebalancing that minimizes edge-cut, which performs equally well but is faster to compute. One explanation is that edge-cut loss and communication cost loss correlate, so moves with small edge-cut loss also incur small communication cost loss. So regardless of the objective function, similar sets of vertices are moved. Additionally, any poor rebalancing move is likely corrected during the label propagation phase due to its high reversal gain.
@@
7:   $\mathcal{M},$ $\Pi^{\prime\prime}\leftarrow\operatorname{UnconLabelProp}(G,\Pi^{\prime})$